A hungry 181 kg lion running northward at 84.5 km/hr attacks and holds onto a 38.6 kg Thomson's gazelle running eastward at 57.4 km/hr. Find the final speed of the lion–gazelle system immediately after the attack. final speed: m/s
Answer #1

Solution)

Applying law of conservation in the north direction,

181 kg* 84.5 km/hr= (181 kg +38.6 kg)*Vy

So, Vy= 69.64 km/hr

Now, applying law of conservation in east direction

38.6 kg*57.4 km/hr= (181 kg + 38.6 kg)* Vx

Vx= 10.08 km/hr

Now,

Velocity, v=sqrt(vx^2 +vy^2)=sqrt(10.08^2 + 69.64^2)= 70.36 km/hr

Direction, tan theta= 69.64/10.08

Theta= 81.76 degrees north of east

=======